# Plasma Instabilities Dominate Radioactive Transients Magnetic Fields: The self-confinement of leptons in Type Ia and Core-Collapse Supernovae, and Kilonovae

### v1

@@ 4.4.1 Overview and Physical Properties
Stripped-Envelope Supernovae (SE SNe) are a subclass of core-collapse supernovae originating from progenitors that have lost their outer hydrogen and, in some cases, helium envelopes before explosion (Clocchiatti et al., 1996; Matheson et al., 2001). This loss of material can occur due to strong stellar winds or binary interactions (e.g., Podsiadlowski et al., 1993; Woosley et al., 1995; Wellstein & Langer, 1999; Wellstein et al., 2001; Heger et al., 2003; Podsiadlowski et al., 2004; Pauldrach et al., 2012; Benvenuto et al., 2013). SE SNe are further categorized based on the presence or absence of hydrogen and helium in their spectra: Type IIb SNe show transient hydrogen features that diminish over time, Type Ib SNe lack hydrogen but show strong helium lines, and Type Ic SNe are devoid of both hydrogen and helium lines, indicating more extensive stripping (Filippenko, 1997).
The progenitors of SE SNe are believed to be massive stars that have undergone significant mass loss. For SNe IIb, the progenitors are often identified as blue supergiants (BSGs), yellow supergiants (YSGs), or red supergiants (RSGs), which retain a thin hydrogen envelope at the time of explosion (Aldering et al., 1994; Crockett et al., 2008; Maund et al., 2004, 2011; Van Dyk et al., 2011, 2014; Smartt, 2015). In contrast, SNe Ib and Ic are believed to originate from Wolf-Rayet (WR) stars, which have strong stellar winds capable of removing both hydrogen and helium layers, or interacting binaries (e.g., Dessart et al., 2011; Georgy et al., 2012; Smartt, 2015; Yoon, 2017). However, there is some debate on the viability of WR stars as progenitors of SE SNe due to limited direct observational evidence (see Yoon et al., 2012). The pre-explosion radii of SE SNe progenitors vary significantly, ranging from about $1$ to $15\,\mathrm{R_{\odot}}$ for WR stars (e.g., Modjaz et al., 2009; Gräfener et al., 2012) and extending up to several hundred solar radii for more extended supergiants. In our analysis, we adopt a typical radius of $15\,\mathrm{R_{\odot}}$, representing a larger WR star, and a range of $1$ to $500\,\mathrm{R_{\odot}}$. Similar to SNe II, we assume iron-core radii before collapse between $1000$ to $2000\,\mathrm{km}$ with a typical value of $\sim 1500\,\mathrm{km}$ for the case of centrally concentrated nickel (e.g., Woosley et al., 2002).
@@ 4.5.1 Overview and Physical Properties
Kilonovae (KNe) are electromagnetic transients powered by the merger of binary neutron stars (NS) or a neutron star and a black hole (BH; Li & Paczyński, 1998; Metzger et al., 2010; Roberts et al., 2011; Barnes & Kasen, 2013). Observations of the gravitational wave event GW170817 (Abbott et al., 2017a) have provided compelling evidence of an optical/infrared kilonova, consistent with theoretical predictions for the electromagnetic counterpart of a binary NS merger (e.g., Abbott et al., 2017b; Arcavi et al., 2017; Chornock et al., 2017; Coulter et al., 2017; Cowperthwaite et al., 2017; Drout et al., 2017; Kasen et al., 2017; Kasliwal et al., 2017; Kilpatrick et al., 2017; McCully et al., 2017; Nicholl et al., 2017; Shappee et al., 2017a; Smartt et al., 2017; Soares-Santos et al., 2017; Tanaka et al., 2017; Tanvir et al., 2017; Piro & Kollmeier, 2018). NS radii, including magnetic and rotating NSs, typically range from $6-20\,\mathrm{km}$, with a common value around $12\,\mathrm{km}$ (e.g., Lattimer & Prakash, 2004; Steiner et al., 2013).
@@ 4.5.2 Magnetic Fields
A comparison between plasma-generated magnetic fields (dark cyan band) and progenitor fields (violet band) is shown in Figure 8d. While progenitor fields decay as $B_{\mathrm{prog}}\propto t^{-2}$ due to the assumption of homologous expansion, starting from values between $10^{7}$ and $10^{15}\,\mathrm{G}$ with a typical value of $10^{12}\,\mathrm{G}$, the instantaneous saturated magnetic fields follow an almost pure power law, approximately $B_{\mathrm{plas}}\propto t^{-1.5}$. The shallower slope of plasma-generated fields ensures they remain stronger than progenitor fields throughout the kilonova’s evolution. This dominance highlights the importance of plasma instabilities in shaping the magnetic field environment of KNe.
@@ 5 Conclusion and Summary
On the observational front, our results open up new pathways for studying the influence of magnetic fields in radioactive transients. Prompt synchrotron emission from trapped positrons, driven by interactions with self-generated magnetic fields, could serve as a direct observational signature of plasma instabilities. Under our highly optimistic scenario, nearby extragalactic SNe Ia, like SN 2011fe and SN 2014J, are unlikely to be detected by LOFAR, SKAO, and ngVLA. For most SNe, CSM interaction will likely dominate the synchrotron emission. Therefore, only a Galactic or local-group SN Ia in a very clean environment may be detectable with LOFAR and SKAO. Further theoretical work, luck (Galactic SN with no CSM), and future generations of radio telescopes are needed to realize the potential of this diagnostic to directly probe the magnetic fields from plasma instabilities in SN ejecta.
While this study establishes the relevance of plasma streaming instabilities, several simplifying assumptions were made, which may impact the first-order details of the results. These assumptions come from numerous uncertainties in both the fields and particles of the systems. First, we assumed that leptons escape as a collimated beam into an unmagnetized environment and are instantaneously injected, which oversimplifies the complexity of ejecta conditions. Nonetheless, the saturated magnetic field should still depend on the free energy of the drifting leptons. Pre-existing magnetic fields, if present, would alter the instability’s specifics without negating its self-confining nature. Second, while the instability growth rate consistently exceeds collisional relaxation times, collisions–neglected in this work–would influence the energy partition between fields and particles at extremely early times. Third, our PIC simulations employed a reduced ion-to-electron mass ratio, which likely affects the final energy distribution, but not the instability’s growth or initial saturation. Additional uncertainties arise from the assumed initial ejecta temperature and ionic composition, as well as the neglect of wave damping, which may influence the long-term evolution of electromagnetic fields. Moreover, the structure of the magnetic power spectrum in $k$-space, critical for lepton escape rates and energy transfer, depends on the details of the initial leptonic distribution function, which warrants further investigation. Future work should incorporate transport models for lepton escape based on diffusion coefficients derived from self-generated fields.
@@ Acknowledgments
The particle-in-cell code tristan-mp v2 is publicly available and the simulated data underlying this paper can be shared upon reasonable request of the corresponding authors.
@@ Appendix B Prompt Synchrotron Emission from SNe Ia Positrons
Relativistic leptons produce synchrotron emission as they interact with magnetic fields. Previous studies have searched for synchrotron emission in nearby SNe Ia from ejecta interaction with circumstellar material (CSM; e.g., Chomiuk et al., 2016; Harris et al., 2023). Taking the assumptions directly from Chomiuk et al. (2016) and extending them to lower frequencies, in Figure 10 top panel, we plot the modeled synchrotron emission from the shocked CSM in SNe Ia for a modest CSM density of $n=0.1\,\mathrm{cm^{-3}}$ in a constant density medium with a fiducial $10\%$ post-shock energy density in magnetic fields.
@@
Figure 10: Top: Synchrotron emission spectral luminosity ($L_{\nu}$) from SNe Ia positrons as a function of frequency ($\nu$) for different times (solid lines) and synchrotron emission from shocked CSM using lower-frequency extension of models from Chomiuk et al. (dashed lines; 2016). Overlaid are the sensitivity curves for ngVLA (dotted line), SKAO (dash-dotted line), and LOFAR (dashed line), assuming 10 hours of integration time. Sensitivity curves are shown for two example distances in different colors: a local-group SN at $1\,\mathrm{Mpc}$ and a Galactic SN at $10\,\mathrm{kpc}$. Bottom: Optical depth ($\tau_{\nu}$) as a function of frequency ($\nu$) at different times from synchrotron self-absorption (SSA; solid lines) and free-free absorption (dashed lines). The color bar indicates the time in days. The cyan dash-dotted line shows where $\tau_{\nu}=1$.
@@
To assess detectability of SNe Ia, we show the sensitivity curves for the next-generation Very Large Array (ngVLA; $1.2-116\,\mathrm{GHz}$; Murphy et al., 2018)555ngVLA sensitivity data was gathered from:
@@
, Square Kilometre Array Observatory (SKAO; $50\,\mathrm{MHz}-50\,\mathrm{GHz}$; Braun et al., 2019) [^6] SKAO sensitivity data was gathered from:
@@
, and the LOw Frequency ARray (LOFAR; van Haarlem et al., 2013), with low-band ($10-80\,\mathrm{MHz}$) and high-band ($110-250\,\mathrm{MHz}$) [^7] LOFAR sensitivity data was gathered from:
@@
It is important to reiterate that the emission results are based on specific simplifying assumptions and optimistic scenario, as detailed above. Although these results do not fully evaluate the detectability of nearby Type Ia SNe, they nonetheless underscore the importance of future work to model the expected radio emission from SNe Ia, particularly at frequencies relevant to future radio observatories. This modeling offers a promising avenue for constraining their synchrotron emission mechanisms.
@@ Appendix C Free neutron decay in kilonovae
Metzger et al. (2015) shows that $\beta$-decay of free neutrons that avoid capture in the outermost layers of KN ejecta power a precursor to the main KN emission on a time scale of hours. Here, we explore the generation of magnetic fields through plasma instabilities in this neutron-rich ejecta, focusing on a simplified model with free neutron decay as the sole energy source.

### v2

@@ 4.4.1 Overview and Physical Properties
SE SNe are a subclass of CC SNe originating from progenitors that have lost their outer hydrogen and, in some cases, helium envelopes before explosion (Clocchiatti et al., 1996; Matheson et al., 2001). This loss of material can occur due to strong stellar winds or binary interactions (e.g., Podsiadlowski et al., 1993; Woosley et al., 1995; Wellstein & Langer, 1999; Wellstein et al., 2001; Heger et al., 2003; Podsiadlowski et al., 2004; Pauldrach et al., 2012; Benvenuto et al., 2013). SE SNe are further categorized based on the presence or absence of hydrogen and helium in their spectra: Type IIb SNe show transient hydrogen features that diminish over time, Type Ib SNe lack hydrogen but show strong helium lines, and Type Ic SNe are devoid of both hydrogen and helium lines, indicating more extensive stripping (Filippenko, 1997).
The progenitors of SE SNe are believed to be massive stars that have undergone significant mass loss. For SNe IIb, the progenitors are often identified as blue supergiants (BSGs), yellow supergiants (YSGs), or RSGs, which retain a thin hydrogen envelope at the time of explosion (Aldering et al., 1994; Crockett et al., 2008; Maund et al., 2004, 2011; Van Dyk et al., 2011, 2014; Smartt, 2015). In contrast, SNe Ib and Ic are believed to originate from Wolf-Rayet (WR) stars, which have strong stellar winds capable of removing both hydrogen and helium layers, or interacting binaries (e.g., Dessart et al., 2011; Georgy et al., 2012; Smartt, 2015; Yoon, 2017). However, there is some debate on the viability of WR stars as progenitors of SE SNe due to limited direct observational evidence (see Yoon et al., 2012). The pre-explosion radii of SE SNe progenitors vary significantly, ranging from about $1$ to $15\,\mathrm{R_{\odot}}$ for WR stars (e.g., Modjaz et al., 2009; Gräfener et al., 2012) and extending up to several hundred solar radii for more extended supergiants. In our analysis, we adopt a typical radius of $15\,\mathrm{R_{\odot}}$, representing a larger WR star, and a range of $1$ to $500\,\mathrm{R_{\odot}}$. Similar to SNe II, we assume iron-core radii before collapse between $1000$ to $2000\,\mathrm{km}$ with a typical value of $\sim 1500\,\mathrm{km}$ for the case of centrally concentrated nickel (e.g., Woosley et al., 2002).
@@ 4.5.1 Overview and Physical Properties
KNe are electromagnetic transients powered by the merger of binary NS or an NS and a black hole (BH; Li & Paczyński, 1998; Metzger et al., 2010; Roberts et al., 2011; Barnes & Kasen, 2013). Observations of the gravitational wave event GW170817 (Abbott et al., 2017a) have provided compelling evidence of an optical/infrared KN, consistent with theoretical predictions for the electromagnetic counterpart of a binary NS merger (e.g., Abbott et al., 2017b; Arcavi et al., 2017; Chornock et al., 2017; Coulter et al., 2017; Cowperthwaite et al., 2017; Drout et al., 2017; Kasen et al., 2017; Kasliwal et al., 2017; Kilpatrick et al., 2017; McCully et al., 2017; Nicholl et al., 2017; Shappee et al., 2017a; Smartt et al., 2017; Soares-Santos et al., 2017; Tanaka et al., 2017; Tanvir et al., 2017; Piro & Kollmeier, 2018). NS radii, including magnetic and rotating NSs, typically range from $6-20\,\mathrm{km}$, with a common value around $12\,\mathrm{km}$ (e.g., Lattimer & Prakash, 2004; Steiner et al., 2013).
@@ 4.5.2 Magnetic Fields
A comparison between plasma-generated magnetic fields (dark cyan band) and progenitor fields (violet band) is shown in Figure 8d. While progenitor fields decay as $B_{\mathrm{prog}}\propto t^{-2}$ due to the assumption of homologous expansion, starting from values between $10^{7}$ and $10^{15}\,\mathrm{G}$ with a typical value of $10^{12}\,\mathrm{G}$, the instantaneous saturated magnetic fields follow an almost pure power law, approximately $B_{\mathrm{plas}}\propto t^{-1.5}$. The shallower slope of plasma-generated fields ensures they remain stronger than progenitor fields throughout the KN’s evolution. This dominance highlights the importance of plasma instabilities in shaping the magnetic field environment of KNe.
@@ 5 Conclusion and Summary
On the observational front, our results open up new pathways for studying the influence of magnetic fields in radioactive transients. Prompt synchrotron emission from trapped positrons, driven by interactions with self-generated magnetic fields, could serve as a direct observational signature of plasma instabilities. Under our highly optimistic scenario, nearby extragalactic SNe Ia, like SN 2011fe and SN 2014J, are unlikely to be detected by LOFAR, SKAO, and ngVLA (next-generation Very Large Array). For most SNe, CSM interaction will likely dominate the synchrotron emission. Therefore, only a Galactic or local-group SN Ia in a very clean environment may be detectable with LOFAR and SKAO. Further theoretical work, luck (Galactic SN with no CSM), and future generations of radio telescopes are needed to realize the potential of this diagnostic to directly probe the magnetic fields from plasma instabilities in SN ejecta.
While this study establishes the relevance of plasma streaming instabilities, several simplifying assumptions were made, which may impact the first-order details of the results. These assumptions come from numerous uncertainties in both the fields and particles of the systems. First, we assumed that leptons escape as a collimated beam into an unmagnetized environment and are instantaneously injected, which oversimplifies the complexity of ejecta conditions. Nonetheless, the saturated magnetic field should still depend on the free energy of the drifting leptons. Furthermore, mechanisms such as continuous particle injection are expected to slightly enhance the magnetic field saturation level and represent an important avenue for future investigation. Pre-existing magnetic fields, if present, would alter the instability’s specifics without negating its self-confining nature. Second, while the instability growth rate consistently exceeds collisional relaxation times, collisions–neglected in this work–would influence the energy partition between fields and particles at extremely early times. Third, our PIC simulations employed a reduced ion-to-electron mass ratio, which likely affects the final energy distribution, but not the instability’s growth or initial saturation.
Additional uncertainties arise from the assumed initial ejecta temperature and ionic composition, as well as the neglect of wave damping, which may influence the long-term evolution of electromagnetic fields. Moreover, the structure of the magnetic power spectrum in $k$-space, critical for lepton escape rates and energy transfer, depends on the details of the initial leptonic distribution function, which warrants further investigation. Future work should incorporate transport models for lepton escape based on diffusion coefficients derived from self-generated fields.
@@ Acknowledgments
AB acknowledges support by the Ministerio de Economía y Competitividad of Spain (Grant No. PID2021-125550OB-I00).
@@
The PIC code tristan-mp v2 is publicly available and the simulated data underlying this paper can be shared upon reasonable request of the corresponding authors.
@@ Appendix B Prompt Synchrotron Emission from SNe Ia Positrons
Relativistic leptons produce synchrotron emission as they interact with magnetic fields. Previous studies have searched for synchrotron emission in nearby SNe Ia from ejecta interaction with CSM (e.g., Chomiuk et al., 2016; Harris et al., 2023). Taking the assumptions directly from Chomiuk et al. (2016) and extending them to lower frequencies, in Figure 10 top panel, we plot the modeled synchrotron emission from the shocked CSM in SNe Ia for a modest CSM density of $n=0.1\,\mathrm{cm^{-3}}$ in a constant density medium with a fiducial $10\%$ post-shock energy density in magnetic fields.
@@
Figure 10: Top: Synchrotron emission spectral luminosity ($L_{\nu}$) from SNe Ia positrons as a function of frequency ($\nu$) for different times (solid lines) and synchrotron emission from shocked CSM using lower-frequency extension of models from Chomiuk et al. (dashed lines; 2016). Overlaid are the sensitivity curves for ngVLA (dotted line), SKAO (dash-dotted line), and LOFAR (dashed line), assuming 10 hours of integration time. Sensitivity curves are shown for two example distances in different colors: a local-group SN at $1\,\mathrm{Mpc}$ and a Galactic SN at $10\,\mathrm{kpc}$. Bottom: Optical depth ($\tau_{\nu}$) as a function of frequency ($\nu$) at different times from SSA (solid lines) and free-free absorption (dashed lines). The color bar indicates the time in days. The cyan dash-dotted line shows where $\tau_{\nu}=1$.
@@
To assess detectability of SNe Ia, we show the sensitivity curves for ngVLA ($1.2-116\,\mathrm{GHz}$; Murphy et al., 2018)555ngVLA sensitivity data were gathered from:
@@
, SKAO ($50\,\mathrm{MHz}-50\,\mathrm{GHz}$; Braun et al., 2019) [^6] SKAO sensitivity data were gathered from:
@@
, and the LOFAR (van Haarlem et al., 2013), with low-band ($10-80\,\mathrm{MHz}$) and high-band ($110-250\,\mathrm{MHz}$) [^7] LOFAR sensitivity data were gathered from:
@@
It is important to reiterate that the emission results are based on specific simplifying assumptions and optimistic scenario, as detailed above. Although these results do not fully evaluate the detectability of nearby SNe Ia, they nonetheless underscore the importance of future work to model the expected radio emission from SNe Ia, particularly at frequencies relevant to future radio observatories. This modeling offers a promising avenue for constraining their synchrotron emission mechanisms.
@@ Appendix C Free neutron decay in kilonovae
Metzger et al. (2015) show that $\beta$-decay of free neutrons that avoid capture in the outermost layers of KN ejecta power a precursor to the main KN emission on a time scale of hours. Here, we explore the generation of magnetic fields through plasma instabilities in this neutron-rich ejecta, focusing on a simplified model with free neutron decay as the sole energy source.